OpenStudy (anonymous):

the sum of three consecutive even integers is 108. Find the integers?

OpenStudy (jdoe0001):

got a number? gimme any number

OpenStudy (aric200):

x+(x+2)+(x+4)=108 3x+6=108 3x=102 x=34 SO your 'three' numbers would be 34, 36, 38.

OpenStudy (anonymous):

My apologies, the number was 195.

OpenStudy (anonymous):

But thank you, man.

OpenStudy (anonymous):

the sum of three consecutive integers is 108. Find the integers?

OpenStudy (anonymous):

I had the correct answer, but whenever I put it in.. it keeps saying incorrect.

OpenStudy (jdoe0001):

ahemm "the sum of three \(\underline{consecutive}\) even integers"

OpenStudy (aric200):

Oh

OpenStudy (aric200):

Idk why I did that then

OpenStudy (jdoe0001):

well.. actually 34, 36 and 38 are indeed ...

OpenStudy (aric200):

No that was right.

OpenStudy (anonymous):

not even, just the sum of three consecutive integers

OpenStudy (aric200):

Now im hella confussed what numbers were even talking about now.

OpenStudy (anonymous):

for 195, not 108. The sum of three consecutive integers is 195. Find the integers.

OpenStudy (jdoe0001):

what the ?

OpenStudy (aric200):

EVEN OR NOT!

OpenStudy (aric200):

You are all over the place.

OpenStudy (anonymous):

The sum of three consecutive integers is 195. Find the integers.

OpenStudy (jdoe0001):

heheh

OpenStudy (aric200):

-_-

OpenStudy (anonymous):

My apologies...

OpenStudy (aric200):

a+a+1+a+2+a+3+a+4=195, 5a+10=195, 5a=185,a=37,they are; 37,38,39,

OpenStudy (aric200):

I gave up on this problem about 20 minutes ago sorry

OpenStudy (texaschic101):

x = 1st number x + 1 = 2nd number x + 2 = 3rd number x + x + 1 + x + 2 = 195 3x + 3 = 195 3x = 195 - 3 3x = 192 x = 192/3 x = 64 x + 1 = 65 x + 2 = 66 your numbers are : 64,65,66